Show me an example problem of how to find the mg of ascorbic acid in a tablet?

Vitamin C tablet

Ascorbic Acid: 176.12g/mol
Answer #1

Suppose that 20ml of a sample was taken from 100ml of a prepared solution. It has 100mg of ascorbic acid in it.

We know that, nearly 49 ml of iodine is required to change colour containing 20ml of ascorbic acid solution. If the solution was done five times to that of the solution, then, total volume of iodine solution = 49 * 5 = 245 ml

Then, Moleiodine = Massascorbic acid * (1/ 176.12* 245*1000) = 0.1 * (1/ 176.12* 245*1000) = 0.00231 mol

For a standard solution,  Mass of ascorbic acid = Moleiodine * 45 * 176.12 = 91.54 mg

Here, amount of ascorbic acid taken was for 100mg and hence, for 250 mg, the ration is (250/100) .